# Towards Modality Generalization: A Benchmark and Prospective Analysis

### v2

@@ 5 Discussions and Outlook
Model selection development. As analyzed before, model selection is important in modality generalization. Without effective model selection, fair benchmarking and reliable performance are difficult to achieve. Future work can focus on (1) task-specific model selection methods: develop and advance techniques tailored to specific MG tasks, such as leave-one-modality-out cross-validation for evaluating generalization; (2) Generalized selection criteria: design selection strategies that rely on surrogate validation sets or meta-learning approaches to ensure robust performance across diverse MG tasks; (3) Adaptive hyperparameter tuning: explore algorithms that adaptively updates hyperparameters based on observed trends during training, reducing dependence on predefined validation data.
@@ Appendix A Implementation Details
Table 3: Details of the learner $h$ with the output dimension of perceptors as 1024 and the number of classes as 20, 23, or 310.
@@
1 Featurizer Linear $1024\rightarrow 512$
@@ 6 ReLU -
7 Linear $512\rightarrow 1024$
8 Classifier (Linear) $1024\rightarrow 20/23/310$
Table 4: Details of hyperparameter setting for random search. That $\mathcal{U}$ denotes the uniform distribution.
@@ 8 Classifier (Linear) $1024\rightarrow 20/23/310$
batchsize $32$ $2^{\mathcal{U}}(3,5.5)$
MML All lr $0.001$ $10^{\mathcal{U}}(-4.5,-2.5)$
momentum $0.9$ $\mathcal{U}(0.85,0.95)$
weight decay $0.0001$ $10^{\mathcal{U}}(-3.5,-4.5)$
patience $70$ $\mathcal{U}(60,80)$
OGM alpha $0.1$ $\mathcal{U}(0.1,0.3)$
DG All lr $0.00005$ $10^{\mathcal{U}}(-5,-3.5)$
weight decay 00 00
IRM lambda $100$ $10^{\mathcal{U}}(-1,5)$
iterations of penalty annealing $500$ $10^{\mathcal{U}}(0,4)$
Mixup alpha $0.2$ $10^{\mathcal{U}}(-1,1)$
CDANN lambda $1.0$ $10^{\mathcal{U}}(-2,-2)$
discriminator weight decay 00 $10^{\mathcal{U}}(-6,-2)$
discriminator steps $1$ $2^{\mathcal{U}}(0,3)$
gradient penalty 00 $10^{\mathcal{U}}(-2,1)$
adam beta1 $0.5$$\{0,0.5\}$
SagNet weight of adversarial loss $0.1$ $10^{\mathcal{U}}(-2,1)$
IB_ERM lambda $100$ $10^{\mathcal{U}}(-1,5)$
iterations of penalty annealing $500$ $10^{\mathcal{U}}(0,4)$
CondCAD lambda $0.1$$\{0.0001,0.001,0.01,0.1,1,10,100\}$
temperature $0.1$$\{0.05,0.1\}$
EQRM lr $0.000001$ $10^{\mathcal{U}}(-7,-5)$
quantile $0.75$ $\mathcal{U}(0.5,0.99)$
iterations of burn-in $2500$ $10^{\mathcal{U}}(2.5,3.5)$
ERM++ lr $0.00005$ $10^{\mathcal{U}}(-5,-3.5)$
URM lambda $0.1$ $\mathcal{U}(0,0.2)$
Hyperparameters. Following DomainBed [52], we list the hyperparameters, their default values, and distributions for our hyperparameter sweep, while involving multimodal learning methods, in Table 4.
Other training details. To optimize the learner $h$, we use Adam [77] for domain generalization algorithms and SGD [78, 79] for multimodal learning methods, following their official implementations. Moreover, we set the training steps to 5,000 for MSR-VTT and the sampled VGGSound-S dataset, and 10,000 for the larger NYUDv2 dataset.
@@ Appendix B Details of Selected Algorithms
  • OGM: On-the-fly Gradient Modulation [59] dynamically adjusts gradients during training to balance contributions from diverse modalities, improving multi-modal learning stability.
  • DLMG: Dynamically Learning Modality Gap [60] adaptively learns to bridge modality-specific differences, facilitating effective integration of heterogeneous data for robust task outcomes.
  • ERM: Empirical Risk Minimization [61] minimizes average loss across training domains, serving as a baseline for domain generalization tasks.
  • IRM: Invariant Risk Minimization [62] seeks invariant predictors across domains, ensuring robust generalization to unseen domains.
  • Mixup: Inter-domain Mixup [63] blends data across domains via interpolation, enhancing model robustness and generalization.
  • CDANN: Class-conditional DANN [48] extends adversarial domain adaptation by aligning class-conditional distributions, improving cross-domain performance.
  • SagNet: Style Agnostic Networks [64] separate style and content, reducing domain-specific biases for better generalization.
@@
  • CondCAD: Conditional Contrastive Adversarial Domain [66] leverages contrastive learning and adversarial methods to align conditional distributions across domains.
  • EQRM: Empirical Quantile Risk Minimization [67] optimizes risk across quantiles, enhancing robustness to domain shifts.
  • ERM++: Improved Empirical Risk Minimization [68] refines ERM with advanced techniques for better cross-domain generalization.
  • URM: Uniform Risk Minimization [69] targets uniform risk across domains, ensuring consistent performance in varied settings.
@@ Appendix C Details of Model Selection Methods
Test-modality validation set (oracle). We select models based on their accuracy on a validation set derived from the test domain distribution. It permits up to multiple queries during hyperparameter tuning with each query corresponding to one hyperparameter configuration. All models are trained for a fixed number of steps without early stopping, and only the final checkpoint is evaluated. While this method can guide model selection, it assumes access to the test modalities and limits the number of hyperparameter combinations [52].
@@ Appendix D Detailed Experimental Results
Table 5: Mean and standard deviation of classification performance comparison under the Weak MG setting w.r.t. training-modality validation set as model selection.
@@
ImageBind MML Concat $\text{23.3}_{\pm\text{3.0}}$ $\text{16.9}_{\pm\text{1.0}}$ $\text{20.4}_{\pm\text{1.4}}$ 20.2 $\text{23.7}_{\pm\text{1.3}}$ $\text{32.6}_{\pm\text{4.4}}$ $\text{32.8}_{\pm\text{4.2}}$ 29.7 $\text{3.5}_{\pm\text{0.6}}$ $\text{2.9}_{\pm\text{0.4}}$ $\text{3.1}_{\pm\text{0.6}}$ 3.2
@@
LanguageBind MML Concat $\text{19.6}_{\pm\text{0.9}}$ $\text{16.8}_{\pm\text{0.6}}$ $\text{19.8}_{\pm\text{1.8}}$ 18.7 $\text{20.2}_{\pm\text{1.7}}$ $\text{40.8}_{\pm\text{7.7}}$ $\text{51.0}_{\pm\text{12.4}}$ 37.3 $\text{2.4}_{\pm\text{0.2}}$ $\text{2.9}_{\pm\text{0.4}}$ $\text{3.5}_{\pm\text{1.0}}$ 3.0
@@
UniBind MML Concat $\text{22.5}_{\pm\text{2.4}}$ $\text{21.1}_{\pm\text{1.6}}$ $\text{18.8}_{\pm\text{3.2}}$ 20.8 $\text{22.2}_{\pm\text{0.5}}$ $\text{35.8}_{\pm\text{6.0}}$ $\text{45.1}_{\pm\text{9.2}}$ 34.4 $\text{2.7}_{\pm\text{0.3}}$ $\text{3.6}_{\pm\text{0.6}}$ $\text{3.2}_{\pm\text{0.5}}$ 3.2
@@
Table 6: Mean and standard deviation of classification performance comparison under the Weak MG setting w.r.t. leave-one-modality-out cross-validation as model selection.
@@
ImageBind MML Concat $\text{11.1}_{\pm\text{0.0}}$ $\text{11.1}_{\pm\text{0.6}}$ $\text{12.2}_{\pm\text{1.0}}$ 11.5 $\text{23.0}_{\pm\text{0.6}}$ $\text{25.2}_{\pm\text{2.4}}$ $\text{22.4}_{\pm\text{0.0}}$ 23.5 $\text{2.3}_{\pm\text{0.0}}$ $\text{1.9}_{\pm\text{0.2}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.2
@@
LanguageBind MML Concat $\text{13.6}_{\pm\text{2.0}}$ $\text{11.8}_{\pm\text{0.1}}$ $\text{14.8}_{\pm\text{1.6}}$ 13.4 $\text{22.3}_{\pm\text{0.0}}$ $\text{22.3}_{\pm\text{0.0}}$ $\text{16.6}_{\pm\text{4.8}}$ 20.4 $\text{2.3}_{\pm\text{0.0}}$ $\text{2.2}_{\pm\text{0.2}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.3
@@
UniBind MML Concat $\text{11.1}_{\pm\text{0.0}}$ $\text{11.1}_{\pm\text{0.6}}$ $\text{11.0}_{\pm\text{0.0}}$ 11.1 $\text{22.3}_{\pm\text{0.0}}$ $\text{31.0}_{\pm\text{5.3}}$ $\text{22.4}_{\pm\text{0.0}}$ 25.2 $\text{2.1}_{\pm\text{0.2}}$ $\text{2.1}_{\pm\text{0.1}}$ $\text{2.3}_{\pm\text{0.1}}$ 2.1
@@
Table 7: Mean and standard deviation of classification performance comparison under the Weak MG setting w.r.t. test-modality validation set (oracle) as model selection.
@@
ImageBind MML Concat $\text{19.1}_{\pm\text{4.0}}$ $\text{15.6}_{\pm\text{1.8}}$ $\text{20.4}_{\pm\text{1.4}}$ 18.4 $\text{22.8}_{\pm\text{0.4}}$ $\text{27.3}_{\pm\text{2.6}}$ $\text{44.3}_{\pm\text{8.9}}$ 31.5 $\text{3.0}_{\pm\text{0.5}}$ $\text{2.9}_{\pm\text{0.3}}$ $\text{3.0}_{\pm\text{0.6}}$ 3.0
@@
LanguageBind MML Concat $\text{16.4}_{\pm\text{2.5}}$ $\text{15.7}_{\pm\text{0.6}}$ $\text{17.2}_{\pm\text{2.6}}$ 16.4 $\text{22.3}_{\pm\text{0.0}}$ $\text{31.8}_{\pm\text{4.1}}$ $\text{45.1}_{\pm\text{12.5}}$ 33.1 $\text{3.1}_{\pm\text{0.6}}$ $\text{3.0}_{\pm\text{0.7}}$ $\text{3.6}_{\pm\text{1.0}}$ 3.2
@@
UniBind MML Concat $\text{19.9}_{\pm\text{4.0}}$ $\text{20.0}_{\pm\text{2.0}}$ $\text{17.9}_{\pm\text{2.9}}$ 19.3 $\text{22.3}_{\pm\text{0.1}}$ $\text{29.8}_{\pm\text{6.1}}$ $\text{33.1}_{\pm\text{4.4}}$ 28.4 $\text{2.5}_{\pm\text{0.3}}$ $\text{3.4}_{\pm\text{0.7}}$ $\text{2.5}_{\pm\text{0.2}}$ 2.8
@@
Table 8: Mean and standard deviation of classification performance comparison under the Strong MG setting w.r.t. training-modality validation set as model selection.
@@
ImageBind MML Concat $\text{11.4}_{\pm\text{0.3}}$ $\text{11.5}_{\pm\text{0.3}}$ $\text{11.0}_{\pm\text{0.0}}$ 11.3 $\text{21.9}_{\pm\text{0.2}}$ $\text{20.8}_{\pm\text{0.7}}$ $\text{22.4}_{\pm\text{0.0}}$ 21.7 $\text{2.3}_{\pm\text{0.0}}$ $\text{2.1}_{\pm\text{0.1}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.3
@@
LanguageBind MML Concat $\text{11.1}_{\pm\text{0.0}}$ $\text{11.2}_{\pm\text{0.6}}$ $\text{10.9}_{\pm\text{0.1}}$ 11.1 $\text{20.7}_{\pm\text{0.6}}$ $\text{22.2}_{\pm\text{0.0}}$ $\text{22.4}_{\pm\text{0.0}}$ 21.8 $\text{2.3}_{\pm\text{0.0}}$ $\text{2.4}_{\pm\text{0.0}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.4
@@
UniBind MML Concat $\text{11.4}_{\pm\text{0.1}}$ $\text{11.4}_{\pm\text{0.3}}$ $\text{11.1}_{\pm\text{0.1}}$ 11.3 $\text{22.4}_{\pm\text{0.1}}$ $\text{21.4}_{\pm\text{0.4}}$ $\text{22.4}_{\pm\text{0.0}}$ 22.1 $\text{2.2}_{\pm\text{0.2}}$ $\text{2.4}_{\pm\text{0.1}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.3
@@
Table 9: Mean and standard deviation of classification performance comparison under the Strong MG setting w.r.t. leave-one-modality-out cross-validation as model selection.
@@
ImageBind MML Concat $\text{8.8}_{\pm\text{1.7}}$ $\text{10.1}_{\pm\text{1.4}}$ $\text{7.4}_{\pm\text{1.3}}$ 8.7 $\text{21.0}_{\pm\text{1.0}}$ $\text{22.3}_{\pm\text{0.0}}$ $\text{20.3}_{\pm\text{1.7}}$ 21.2 $\text{2.2}_{\pm\text{0.1}}$ $\text{2.1}_{\pm\text{0.2}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.2
@@
LanguageBind MML Concat $\text{7.0}_{\pm\text{1.7}}$ $\text{11.8}_{\pm\text{0.1}}$ $\text{9.4}_{\pm\text{1.3}}$ 9.4 $\text{22.3}_{\pm\text{0.0}}$ $\text{22.3}_{\pm\text{0.0}}$ $\text{22.4}_{\pm\text{0.0}}$ 22.3 $\text{2.2}_{\pm\text{0.1}}$ $\text{1.4}_{\pm\text{0.5}}$ $\text{2.0}_{\pm\text{0.3}}$ 1.9
@@
UniBind MML Concat $\text{9.9}_{\pm\text{1.0}}$ $\text{8.8}_{\pm\text{1.2}}$ $\text{8.9}_{\pm\text{1.1}}$ 9.2 $\text{14.8}_{\pm\text{3.8}}$ $\text{20.8}_{\pm\text{1.1}}$ $\text{12.3}_{\pm\text{5.5}}$ 16.0 $\text{2.3}_{\pm\text{0.0}}$ $\text{2.2}_{\pm\text{0.2}}$ $\text{2.2}_{\pm\text{0.1}}$ 2.2
@@
Table 10: Mean and standard deviation of classification performance comparison under the Strong MG setting w.r.t. test-modality validation set (oracle) as model selection.
@@
ImageBind MML Concat $\text{11.1}_{\pm\text{0.0}}$ $\text{11.8}_{\pm\text{0.1}}$ $\text{11.0}_{\pm\text{0.0}}$ 11.3 $\text{22.3}_{\pm\text{0.0}}$ $\text{22.3}_{\pm\text{0.0}}$ $\text{22.4}_{\pm\text{0.0}}$ 22.3 $\text{2.3}_{\pm\text{0.0}}$ $\text{2.4}_{\pm\text{0.0}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.4
@@
LanguageBind MML Concat $\text{11.1}_{\pm\text{0.0}}$ $\text{11.8}_{\pm\text{0.1}}$ $\text{11.0}_{\pm\text{0.0}}$ 11.3 $\text{22.3}_{\pm\text{0.0}}$ $\text{22.3}_{\pm\text{0.0}}$ $\text{22.4}_{\pm\text{0.0}}$ 22.3 $\text{2.2}_{\pm\text{0.1}}$ $\text{2.4}_{\pm\text{0.0}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.3

### v3

@@ 7.4. Model Selection Development
Model selection is important in modality generalization. Without effective model selection, fair benchmarking and reliable performance are difficult to achieve. Future work can focus on (1) task-specific model selection: develop and advance techniques tailored to specific MG tasks, such as leave-one-modality-out cross-validation for evaluating generalization; (2) generalized selection criteria: design selection strategies that rely on surrogate validation sets or meta-learning approaches to ensure robust performance across diverse MG tasks; (3) adaptive hyperparameter tuning: explore algorithms that adaptively update hyperparameters based on observed trends during training, reducing dependence on predefined validation data.
7.5. Limitations
This work proposes a comprehensive benchmark for the new scenario of modality generalization. In particular, we investigate five modalities across three datasets. Additional modalities in our daily lives, such as bioelectric signals, are also promising for advancing modality generalization. Unfortunately, we cannot curate these data due to privacy concerns and data accessibility issues. These types of data often involve sensitive personal or medical information, requiring strict compliance with ethical and legal regulations. More modalities like thermal imaging, LiDAR scans, and haptic feedback also hold potential significance in pushing modality generalization, especially for robotics and autonomous systems. These data are scarce due to the overhead collection and labeling costs. The lack of corresponding modality encoders further hinders us from including these modalities currently. With community development, we will continue ModalBed, as emphasized in Section 4.
Acknowledgement
This research/project is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP2024-002). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore. Xiaobo Xia is supported by MoE Key Laboratory of Brain-inspired Intelligent Perception and Cognition, University of Science and Technology of China (Grant No. 2421002).
@@ Appendix A Implementation Details
Table 3. Details of the learner $h$ with the output dimension of perceptors as 1024 and the number of classes as 20, 23, or 310.
@@
1
Featurizer
Linear 10245121024$\rightarrow$ 5121024 → 512
@@ 6 ReLU -
7 Linear 5121024512$\rightarrow$ 1024512 → 1024
8 Classifier (Linear) 102420/23/3101024$\rightarrow$ 20/23/3101024 → 20 / 23 / 310
Table 4. Details of hyperparameter setting for random search. That $\mathcal{U}$ denotes the uniform distribution.
@@ 8 Classifier (Linear) 102420/23/3101024$\rightarrow$ 20/23/3101024 → 20 / 23 / 310
batchsize 323232 $2^{\mathcal{U}}(3,5.5)$
MML All lr 0.0010.0010.001 $10^{\mathcal{U}}(-4.5,-2.5)$
momentum 0.90.90.9 $\mathcal{U}(0.85,0.95)$
weight decay 0.00010.00010.0001 $10^{\mathcal{U}}(-3.5,-4.5)$
patience 707070 $\mathcal{U}(60,80)$
OGM alpha 0.10.10.1 $\mathcal{U}(0.1,0.3)$
DG All lr 0.000050.000050.00005 $10^{\mathcal{U}}(-5,-3.5)$
weight decay 0 0
IRM lambda 100100100 $10^{\mathcal{U}}(-1,5)$
iterations of penalty annealing 500500500 $10^{\mathcal{U}}(0,4)$
Mixup alpha 0.20.20.2 $10^{\mathcal{U}}(-1,1)$
CDANN lambda 1.01.01.0 $10^{\mathcal{U}}(-2,-2)$
discriminator weight decay 0 $10^{\mathcal{U}}(-6,-2)$
discriminator steps 111 $2^{\mathcal{U}}(0,3)$
gradient penalty 0 $10^{\mathcal{U}}(-2,1)$
[…]
SagNet weight of adversarial loss 0.10.10.1 $10^{\mathcal{U}}(-2,1)$
IB_ERM lambda 100100100 $10^{\mathcal{U}}(-1,5)$
iterations of penalty annealing 500500500 $10^{\mathcal{U}}(0,4)$
[…]
EQRM lr 0.0000010.0000010.000001 $10^{\mathcal{U}}(-7,-5)$
quantile 0.750.750.75 $\mathcal{U}(0.5,0.99)$
iterations of burn-in 250025002500 $10^{\mathcal{U}}(2.5,3.5)$
ERM++ lr 0.000050.000050.00005 $10^{\mathcal{U}}(-5,-3.5)$
URM lambda 0.10.10.1 $\mathcal{U}(0,0.2)$
Hyperparameters. Following DomainBed (Gulrajani and Lopez-Paz, 2021), we list the hyperparameters, their default values, and distributions for our hyperparameter sweep, while involving multimodal learning methods, in Table 4.
Other training details. To optimize the learner $h$, we use Adam (Kingma, 2014) for domain generalization algorithms and SGD (Robbins and Monro, 1951; LeCun et al., 1998) for multimodal learning methods, following their official implementations. Moreover, we set the training steps to 5,000 for MSR-VTT and the sampled VGGSound-S dataset, and 10,000 for the larger NYUDv2 dataset.
@@ Appendix B Details of Selected Algorithms
  • OGM: On-the-fly Gradient Modulation (Peng et al., 2022) dynamically adjusts gradients during training to balance contributions from diverse modalities, improving multi-modal learning stability.
  • DLMG: Dynamically Learning Modality Gap (Yang et al., 2024b) adaptively learns to bridge modality-specific differences, facilitating effective integration of heterogeneous data for robust task outcomes.
  • ERM: Empirical Risk Minimization (Vapnik et al., 1998) minimizes average loss across training domains, serving as a baseline for domain generalization tasks.
  • IRM: Invariant Risk Minimization (Arjovsky et al., 2019) seeks invariant predictors across domains, ensuring robust generalization to unseen domains.
  • Mixup: Inter-domain Mixup (Yan et al., 2020) blends data across domains via interpolation, enhancing model robustness and generalization.
  • CDANN: Class-conditional DANN (Li et al., 2018a) extends adversarial domain adaptation by aligning class-conditional distributions, improving cross-domain performance.
  • SagNet: Style Agnostic Networks (Nam et al., 2021) separate style and content, reducing domain-specific biases for better generalization.
@@
  • CondCAD: Conditional Contrastive Adversarial Domain (Ruan et al., 2022) leverages contrastive learning and adversarial methods to align conditional distributions across domains.
  • EQRM: Empirical Quantile Risk Minimization (Eastwood et al., 2022) optimizes risk across quantiles, enhancing robustness to domain shifts.
  • ERM++: Improved Empirical Risk Minimization (Teterwak et al., 2023) refines ERM with advanced techniques for better cross-domain generalization.
  • URM: Uniform Risk Minimization (Krishnamachari et al., 2024) targets uniform risk across domains, ensuring consistent performance in varied settings.
@@ Appendix C Details of Model Selection Methods
Test-modality validation set (oracle). We select models based on their accuracy on a validation set derived from the test domain distribution. It permits up to multiple queries during hyperparameter tuning with each query corresponding to one hyperparameter configuration. All models are trained for a fixed number of steps without early stopping, and only the final checkpoint is evaluated. While this method can guide model selection, it assumes access to the test modalities and limits the number of hyperparameter combinations (Gulrajani and Lopez-Paz, 2021).
@@ Appendix D Detailed Experimental Results
Table 5. Mean and standard deviation of classification performance comparison under the Weak MG setting w.r.t. training-modality validation set as model selection.
@@
ImageBind
MML Concat $\text{23.3}_{\pm\text{3.0}}$ $\text{16.9}_{\pm\text{1.0}}$ $\text{20.4}_{\pm\text{1.4}}$ 20.2 $\text{23.7}_{\pm\text{1.3}}$ $\text{32.6}_{\pm\text{4.4}}$ $\text{32.8}_{\pm\text{4.2}}$ 29.7 $\text{3.5}_{\pm\text{0.6}}$ $\text{2.9}_{\pm\text{0.4}}$ $\text{3.1}_{\pm\text{0.6}}$ 3.2
@@
LanguageBind
MML Concat $\text{19.6}_{\pm\text{0.9}}$ $\text{16.8}_{\pm\text{0.6}}$ $\text{19.8}_{\pm\text{1.8}}$ 18.7 $\text{20.2}_{\pm\text{1.7}}$ $\text{40.8}_{\pm\text{7.7}}$ $\text{51.0}_{\pm\text{12.4}}$ 37.3 $\text{2.4}_{\pm\text{0.2}}$ $\text{2.9}_{\pm\text{0.4}}$ $\text{3.5}_{\pm\text{1.0}}$ 3.0
@@
UniBind
MML Concat $\text{22.5}_{\pm\text{2.4}}$ $\text{21.1}_{\pm\text{1.6}}$ $\text{18.8}_{\pm\text{3.2}}$ 20.8 $\text{22.2}_{\pm\text{0.5}}$ $\text{35.8}_{\pm\text{6.0}}$ $\text{45.1}_{\pm\text{9.2}}$ 34.4 $\text{2.7}_{\pm\text{0.3}}$ $\text{3.6}_{\pm\text{0.6}}$ $\text{3.2}_{\pm\text{0.5}}$ 3.2
@@
Table 6. Mean and standard deviation of classification performance comparison under the Weak MG setting w.r.t. leave-one-modality-out cross-validation as model selection.
@@
ImageBind
MML Concat $\text{11.1}_{\pm\text{0.0}}$ $\text{11.1}_{\pm\text{0.6}}$ $\text{12.2}_{\pm\text{1.0}}$ 11.5 $\text{23.0}_{\pm\text{0.6}}$ $\text{25.2}_{\pm\text{2.4}}$ $\text{22.4}_{\pm\text{0.0}}$ 23.5 $\text{2.3}_{\pm\text{0.0}}$ $\text{1.9}_{\pm\text{0.2}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.2
@@
LanguageBind
MML Concat $\text{13.6}_{\pm\text{2.0}}$ $\text{11.8}_{\pm\text{0.1}}$ $\text{14.8}_{\pm\text{1.6}}$ 13.4 $\text{22.3}_{\pm\text{0.0}}$ $\text{22.3}_{\pm\text{0.0}}$ $\text{16.6}_{\pm\text{4.8}}$ 20.4 $\text{2.3}_{\pm\text{0.0}}$ $\text{2.2}_{\pm\text{0.2}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.3
@@
UniBind
MML Concat $\text{11.1}_{\pm\text{0.0}}$ $\text{11.1}_{\pm\text{0.6}}$ $\text{11.0}_{\pm\text{0.0}}$ 11.1 $\text{22.3}_{\pm\text{0.0}}$ $\text{31.0}_{\pm\text{5.3}}$ $\text{22.4}_{\pm\text{0.0}}$ 25.2 $\text{2.1}_{\pm\text{0.2}}$ $\text{2.1}_{\pm\text{0.1}}$ $\text{2.3}_{\pm\text{0.1}}$ 2.1
@@
Table 7. Mean and standard deviation of classification performance comparison under the Weak MG setting w.r.t. test-modality validation set (oracle) as model selection.
@@
ImageBind
MML Concat $\text{19.1}_{\pm\text{4.0}}$ $\text{15.6}_{\pm\text{1.8}}$ $\text{20.4}_{\pm\text{1.4}}$ 18.4 $\text{22.8}_{\pm\text{0.4}}$ $\text{27.3}_{\pm\text{2.6}}$ $\text{44.3}_{\pm\text{8.9}}$ 31.5 $\text{3.0}_{\pm\text{0.5}}$ $\text{2.9}_{\pm\text{0.3}}$ $\text{3.0}_{\pm\text{0.6}}$ 3.0
@@
LanguageBind
MML Concat $\text{16.4}_{\pm\text{2.5}}$ $\text{15.7}_{\pm\text{0.6}}$ $\text{17.2}_{\pm\text{2.6}}$ 16.4 $\text{22.3}_{\pm\text{0.0}}$ $\text{31.8}_{\pm\text{4.1}}$ $\text{45.1}_{\pm\text{12.5}}$ 33.1 $\text{3.1}_{\pm\text{0.6}}$ $\text{3.0}_{\pm\text{0.7}}$ $\text{3.6}_{\pm\text{1.0}}$ 3.2
@@
UniBind
MML Concat $\text{19.9}_{\pm\text{4.0}}$ $\text{20.0}_{\pm\text{2.0}}$ $\text{17.9}_{\pm\text{2.9}}$ 19.3 $\text{22.3}_{\pm\text{0.1}}$ $\text{29.8}_{\pm\text{6.1}}$ $\text{33.1}_{\pm\text{4.4}}$ 28.4 $\text{2.5}_{\pm\text{0.3}}$ $\text{3.4}_{\pm\text{0.7}}$ $\text{2.5}_{\pm\text{0.2}}$ 2.8
@@
Table 8. Mean and standard deviation of classification performance comparison under the Strong MG setting w.r.t. training-modality validation set as model selection.
@@
ImageBind
MML Concat $\text{11.4}_{\pm\text{0.3}}$ $\text{11.5}_{\pm\text{0.3}}$ $\text{11.0}_{\pm\text{0.0}}$ 11.3 $\text{21.9}_{\pm\text{0.2}}$ $\text{20.8}_{\pm\text{0.7}}$ $\text{22.4}_{\pm\text{0.0}}$ 21.7 $\text{2.3}_{\pm\text{0.0}}$ $\text{2.1}_{\pm\text{0.1}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.3
@@
LanguageBind
MML Concat $\text{11.1}_{\pm\text{0.0}}$ $\text{11.2}_{\pm\text{0.6}}$ $\text{10.9}_{\pm\text{0.1}}$ 11.1 $\text{20.7}_{\pm\text{0.6}}$ $\text{22.2}_{\pm\text{0.0}}$ $\text{22.4}_{\pm\text{0.0}}$ 21.8 $\text{2.3}_{\pm\text{0.0}}$ $\text{2.4}_{\pm\text{0.0}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.4
@@
UniBind
MML Concat $\text{11.4}_{\pm\text{0.1}}$ $\text{11.4}_{\pm\text{0.3}}$ $\text{11.1}_{\pm\text{0.1}}$ 11.3 $\text{22.4}_{\pm\text{0.1}}$ $\text{21.4}_{\pm\text{0.4}}$ $\text{22.4}_{\pm\text{0.0}}$ 22.1 $\text{2.2}_{\pm\text{0.2}}$ $\text{2.4}_{\pm\text{0.1}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.3
@@
Table 9. Mean and standard deviation of classification performance comparison under the Strong MG setting w.r.t. leave-one-modality-out cross-validation as model selection.
@@
ImageBind
MML Concat $\text{8.8}_{\pm\text{1.7}}$ $\text{10.1}_{\pm\text{1.4}}$ $\text{7.4}_{\pm\text{1.3}}$ 8.7 $\text{21.0}_{\pm\text{1.0}}$ $\text{22.3}_{\pm\text{0.0}}$ $\text{20.3}_{\pm\text{1.7}}$ 21.2 $\text{2.2}_{\pm\text{0.1}}$ $\text{2.1}_{\pm\text{0.2}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.2
@@
LanguageBind
MML Concat $\text{7.0}_{\pm\text{1.7}}$ $\text{11.8}_{\pm\text{0.1}}$ $\text{9.4}_{\pm\text{1.3}}$ 9.4 $\text{22.3}_{\pm\text{0.0}}$ $\text{22.3}_{\pm\text{0.0}}$ $\text{22.4}_{\pm\text{0.0}}$ 22.3 $\text{2.2}_{\pm\text{0.1}}$ $\text{1.4}_{\pm\text{0.5}}$ $\text{2.0}_{\pm\text{0.3}}$ 1.9
@@
UniBind
MML Concat $\text{9.9}_{\pm\text{1.0}}$ $\text{8.8}_{\pm\text{1.2}}$ $\text{8.9}_{\pm\text{1.1}}$ 9.2 $\text{14.8}_{\pm\text{3.8}}$ $\text{20.8}_{\pm\text{1.1}}$ $\text{12.3}_{\pm\text{5.5}}$ 16.0 $\text{2.3}_{\pm\text{0.0}}$ $\text{2.2}_{\pm\text{0.2}}$ $\text{2.2}_{\pm\text{0.1}}$ 2.2
@@
Table 10. Mean and standard deviation of classification performance comparison under the Strong MG setting w.r.t. test-modality validation set (oracle) as model selection.
@@
ImageBind
MML Concat $\text{11.1}_{\pm\text{0.0}}$ $\text{11.8}_{\pm\text{0.1}}$ $\text{11.0}_{\pm\text{0.0}}$ 11.3 $\text{22.3}_{\pm\text{0.0}}$ $\text{22.3}_{\pm\text{0.0}}$ $\text{22.4}_{\pm\text{0.0}}$ 22.3 $\text{2.3}_{\pm\text{0.0}}$ $\text{2.4}_{\pm\text{0.0}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.4
@@
LanguageBind
MML Concat $\text{11.1}_{\pm\text{0.0}}$ $\text{11.8}_{\pm\text{0.1}}$ $\text{11.0}_{\pm\text{0.0}}$ 11.3 $\text{22.3}_{\pm\text{0.0}}$ $\text{22.3}_{\pm\text{0.0}}$ $\text{22.4}_{\pm\text{0.0}}$ 22.3 $\text{2.2}_{\pm\text{0.1}}$ $\text{2.4}_{\pm\text{0.0}}$ $\text{2.3}_{\pm\text{0.0}}$ 2.3